OpenStudy (jusaquikie):

lim 8e^(TanX) x → (π/2)+

OpenStudy (jusaquikie):

not sure where to go to with this

OpenStudy (turingtest):

since\[\Large\lim_{x\to a}e^{f(x)}=e^{\lim_{x\to a}f(x)}\]really all you need to know is the limit\[\Large\lim_{x\to\pi/2^+}\tan x\]

OpenStudy (turingtest):

what is\[\lim_{x\to\pi/2}\tan x\]

OpenStudy (turingtest):

?

OpenStudy (jusaquikie):

+infinity?

OpenStudy (jusaquikie):

.027

OpenStudy (turingtest):

actually it depends on the left and right hand approach from the left (x<pi/2) cos x>0 from the right (x>pi/2) cos x<0

OpenStudy (turingtest):

I have no idea where you got that number from....

OpenStudy (jusaquikie):

tan(pi/2) lol

OpenStudy (turingtest):

I'm gonna take a wild guess and say you left your calculator in degree mode

OpenStudy (jusaquikie):

yes

OpenStudy (turingtest):

\[\tan x=\frac{\sin x}{\cos x}\]so\[\tan (\pi/2)=\frac{\sin (\pi/2)}{\cos (\pi/2)}=?\]

OpenStudy (jusaquikie):

1/0= undefined

OpenStudy (turingtest):

right, now what about coming from the right, x>pi/2 will cosince be positive or negative approaching pi/2 from the right?

OpenStudy (turingtest):

cosine*

OpenStudy (jusaquikie):

positive and increasing to 1?

OpenStudy (turingtest):

no, what is the cosine of pi/2 ?

OpenStudy (jusaquikie):

0 sorry was thinking sin

OpenStudy (jusaquikie):

so negative

OpenStudy (turingtest):

correct, so as \(x\to\pi/2^+\) we have that \(\cos x\to0\), which means that\[\lim_{x\to\pi/2^+}\tan x=?\]

OpenStudy (jusaquikie):

0

OpenStudy (turingtest):

no, think in terms of sin and cos

OpenStudy (jusaquikie):

anything with Euler's number just confuses me, i'm not sure how to treat it

OpenStudy (jusaquikie):

2pi

OpenStudy (turingtest):

ignore Euler's number, it could be any exponential base, the answer would be the same...\[\lim_{x\to\pi/2^+}\frac{\sin x}{\cos x}=?\]what is sine approaching? what is cos approaching?

OpenStudy (turingtest):

what is sine approaching? what is cos approaching?

OpenStudy (jusaquikie):

1,0

OpenStudy (turingtest):

and is that zero being approached from the negative or positive side?

OpenStudy (jusaquikie):

negative?

OpenStudy (turingtest):

correct, so considering that\[\lim_{x\to\pi/2^+}\tan x=\lim_{x\to\pi/2^+}\frac{\sin x}{\cos x}\to\frac10\]and that that zero is being approached from the negative side, what is the limit?

OpenStudy (jusaquikie):

i can't visualize it and i'm not sure how to graph it in my calculator so i'm trying to relate it to the unit circle

OpenStudy (jusaquikie):

-infinity?

OpenStudy (jusaquikie):

no + infinity

OpenStudy (turingtest):

you were right the first time, -infty